James Tung* (yanchun.tung@gmail.com), Department of Mathematics, DePaul University, 2320 N. Kenmore Ave., Chicago, IL 60614. Taylor coefficients of functions in Fock spaces. For analytic functions on the unit disk, the classical Hausdorff-Young and Hardy-Littlewood theorems describe necessary or sufficient conditions for the functions to belong to Hardy spaces. The conditions are in terms of the Taylor coefficients of the functions. We will show that there are analogous conditions for entire functions to belong to Fock spaces, and that these conditions can be improved in certain senses.